AHi. I was wondering if I could recieve help on the following problem:
I know you do order of operations, but I do not know if it matters if I multiply or divide first after I get the exponent.
4 - 4^2 divided by 8 x 3 + 6
What I did was get 4 -16 divided by 8 x 3 + 6
Then if I multiplied first, I would get 4 - 16 divided by 24 + 6
Then I would get 4 - 2/3 + 6. Then I'd get 9 and 1/3.
On my calculator I got 4, but that would only work if I divided before I multiplied. I know division and multiplication are inverse opperations, so would it be okay if I divided first and got 4 as the answer?
Thanks.

Found 2 solutions by stanbon, edjones:
Answer by stanbon(57424)

4 - 4^2 divided by 8 x 3 + 6
-----------
Is this your problem?
4-16/8*3+6
=4-2*3+6
=4-6+6
=4
-----------
OR is this your problem?
[4-4^2]/[8*3+6]
= -12/30
= -2/5
----------------
Comment: When you write "divided by" the reader cannot
tell where the denominator ends. Please use parentheses
so we can see the problem the way you see it.
==============
Comment:
The general rule for evaluating problems of the type
you posted is
1st: multiply and divide from left to right in the order they occur.
2nd: add and subtract from left to right in the order they occur.
================
Cheers,
Stan H.

Answer by edjones(7569)

Tip #1 don't use x for a times sign. Use *
then you wont get the x's in your problems mixed up with your times signs.
Tip #2 use brackets ( ) or [ ] etc. to avoid confusion. Eg: 8*3+6 could be 8(3+6).
Whether you do multipliction or division 1st doesn't make any difference but the way you set up the problem does.
If you want to divide 1st (4-16)/((8*3)+6)) is the way to set it up.
the calculator will give -0.4 as the answer.
With pencil and paper we do the operations inside the brackets 1st.
(4-16)=-12
(8*3)=24
(24+6)=30
-12/30=-0.4
EdJones